OpenStudy (anonymous):

integrate sqrt(16-y^2)/2 dy

OpenStudy (shamil98):

\[\large \sqrt{16-y^2} = (16-y^2)^\frac{ 1 }{ 2 }\]

OpenStudy (shamil98):

There isn't a rule for integration like the quotient rule is there?..

OpenStudy (anonymous):

integration by parts or integration by subtitution

OpenStudy (shamil98):

yeah, i'm looking at that right now.

OpenStudy (shamil98):

First choose u and v u = (16-y^2)^1/2 v = 1/2

OpenStudy (shamil98):

differentiate u and integrate v.

OpenStudy (shamil98):

ya with me so far?

OpenStudy (shamil98):

we're using the formula \[\large \int\limits_{}^{} udv = uv - \int\limits_{}^{} v~du\]

OpenStudy (anonymous):

erm, wait me read up for a while >< coz preparing test for other subject tmr, n doing my engineering math assignment simultaneously

OpenStudy (shamil98):

use the chain rule for u and integrate v (16-y^2)^1/2 = u u' = -2y/(16-y^2)^1/2 v = 1/2 integrating v = x/2 + c

OpenStudy (shamil98):

I'm just following an example lol , not sure if i'm doing this correctly >.<

OpenStudy (anonymous):

ok, nvm, we can study together coz i left my integration by parts and integration by substitution abt 1 year ald

OpenStudy (anonymous):

i get the fundamental of integration by part ald